# Safety Controller Synthesis for Stochastic Networked Systems under Communication Constraints

### v1

@@ II-C Augmented dt-SLS-$\Sigma$
To formally analyze the safety of the networked control system dt-SLS-$\Sigma$, we first introduce an augmented networked system representation that captures both the inputs and states of the dt-SLS-$\Sigma$ and serves as the foundation for defining the CBC, which follows next. The dynamics of $x_{k}$ and $\hat{x}_{k}^{c}$ are given in (3) and (4), while dynamics of $u_{k}$ and $\hat{u}_{k}^{c}$ can be defined as
@@
where $\mathcal{A}\in\mathbb{R}^{\kappa\times\kappa}$ and $\mathcal{D}\in\mathbb{R}^{\kappa\times\psi}$. One can write (10) as
$\!\!\Lambda\!$ $:\!\begin{bmatrix}{\mathbf{x}}_{k+1}\\ \hat{\mathbf{x}}_{k+1}^{c}\\ {\mathbf{u}}_{k+1}\\ \hat{\mathbf{u}}_{k+1}^{c}\end{bmatrix}\!=\!\underbrace{\begin{bmatrix}\mathcal{A}_{11}&\mathcal{A}_{12}&\mathcal{A}_{13}&\mathcal{A}_{14}\\ \mathcal{A}_{21}&\mathcal{A}_{22}&\mathcal{A}_{23}&\mathcal{A}_{24}\\ \mathcal{A}_{31}&\mathcal{A}_{32}&\mathcal{A}_{33}&\mathcal{A}_{34}\\ \mathcal{A}_{41}&\mathcal{A}_{42}&\mathcal{A}_{43}&\mathcal{A}_{44}\end{bmatrix}}_{\mathcal{A}}\!\begin{bmatrix}{\mathbf{x}}_{k}\\ \hat{\mathbf{x}}_{k}^{c}\\ {\mathbf{u}}_{k}\\ \hat{\mathbf{u}}_{k}^{c}\end{bmatrix}\!+\mathcal{D}\mathbf{w}_{k},$
$\text{with}~{}\mathcal{D}$ $=\begin{bmatrix}\mathbf{I}_{n}&{A}&\dots&{A}^{\tau}\\ \boldsymbol{0}_{(\kappa-n)\times n}&\boldsymbol{0}_{(\kappa-n)\times n}&\dots&\boldsymbol{0}_{(\kappa-n)\times n}\end{bmatrix}_{\kappa\times\psi}\!\!,$ (11)
and $\mathcal{A}_{ij},~{}i,j\in\{1,\dots,4\}$ are defined as in (23) in the Appendix. Since $\phi_{k-t-1}$ is an i.i.d. Bernoulli random variable for all $t\in\{0,1,\dots,\tau-1\}$, we simplify the notation by denoting $\phi_{k-t-1}$ as $\phi$ and $1-\phi_{k-t-1}$ as $\bar{\phi}$. We used the padding technique [24] in constructing matrix $\mathcal{A}$ to explicitly represent the state variable transitions from one time instant to the next. The padding incorporates trivial equations $x_{i}=x_{i}$, with $\forall i\in[k-\tau+1,k]$ and $\hat{x}_{k}^{c}=\hat{x}_{k}^{c}$ ensuring continuity in the state representation of the augmented system dt-ASLS-$\Lambda$.
We now decompose matrix $\mathcal{A}$ in (11) into a sum of multiple components. Specifically, $\mathcal{A}$ is expressed as a combination of a baseline matrix $\mathcal{A}_{1}$ and additional terms that depend on the Bernoulli-distributed variable $\phi$. This decomposition allows us to analyze the dynamics of the augmented system dt-SLS-$\Lambda$ by isolating the effect of $\phi$. Furthermore, we rewrite the Bernoulli variable $\phi=q_{\phi}(1-\zeta_{\phi})$, where the new variable $\zeta_{\phi}$ has a two-point distribution and takes one of two possible values: either $1$ or $1-\frac{1}{q_{\phi}}$, with $\mathds{E}[\zeta_{\phi}]=0$ and $\mathsf{Var}[\zeta_{\phi}]=\mathds{E}[\zeta_{\phi}^{2}]=\frac{1-q_{\phi}}{q_{\phi}}$. This transformation simplifies the computation by allowing the decomposition of $\mathcal{A}$ as
@@
$\mathcal{A}_{1}$ $\!\!=\!\!\begin{bmatrix}\mathcal{A}_{11}\!\!&\!\!\mathcal{A}_{12}\!\!&\!\!\mathcal{A}_{13}^{a}\!\!&\!\!\mathcal{A}_{14}^{a}\\ {\mathcal{A}}_{21}\!\!&\!\!{\mathcal{A}}_{22}\!\!&\!\!\mathcal{A}_{23}\!\!&\!\!\mathcal{A}_{24}\\ \!\mathbf{0}_{(\varpi+m)\times\psi}\!\!&\!\!\mathbf{0}_{(\varpi+m)\times 2n}\!\!&\!\!\mathbf{0}_{(\varpi+m)\times(\varpi+m)}\!\!&\!\!\mathbf{0}_{(\varpi+m)\times\varpi}\!\!\!\\ \mathbf{0}_{\varpi\times\psi}\!\!&\!\!\mathbf{0}_{\varpi\times 2n}\!\!&\!\!\mathbf{0}_{\varpi\times(\varpi+m)}\!\!&\!\!\mathbf{0}_{\varpi\times\varpi}\end{bmatrix}\!\!,$
$\mathcal{A}_{2}$ $\!\!=\!\!\begin{bmatrix}\mathbf{0}_{\psi\times\psi}\!&\!\mathbf{0}_{\psi\times 2n}\!&\!q_{\phi}\mathcal{A}_{13}^{b}\!&\!(1-q_{\phi})\mathcal{A}_{14}^{b}\\ \mathbf{0}_{2n\times\psi}\!&\!\mathbf{0}_{2n\times 2n}\!&\!\mathbf{0}_{2n\times(\varpi+m)}\!&\!\mathbf{0}_{2n\times\varpi}\\ q_{\phi}\mathcal{A}_{31}^{a}\!&\!q_{\phi}\mathcal{A}_{32}^{a}\!&\!(1-q_{\phi})\mathcal{A}_{33}^{a}\!&\!q_{\phi}\mathcal{A}_{34}^{a}\\ \mathbf{0}_{\varpi\times\psi}\!&\!\mathbf{0}_{\varpi\times 2n}\!&\!\mathbf{0}_{\varpi\times(\varpi+m)}\!&\!\mathbf{0}_{\varpi\times\varpi}\end{bmatrix}\!\!,$
$\mathcal{A}_{3}$ $\!\!=\!\!\begin{bmatrix}\mathbf{0}_{\psi\times\psi}\!&\!\mathbf{0}_{\psi\times 2n}\!&\!-q_{\phi}\mathcal{A}_{13}^{b}\!&\!q_{\phi}\mathcal{A}_{14}^{b}\\ \mathbf{0}_{2n\times\psi}\!&\!\mathbf{0}_{2n\times 2n}\!&\!\mathbf{0}_{2n\times(\varpi+m)}\!&\!\mathbf{0}_{2n\times\varpi}\\ -q_{\phi}\mathcal{A}_{31}^{a}\!&\!-q_{\phi}\mathcal{A}_{32}^{a}\!&\!q_{\phi}\mathcal{A}_{33}^{a}\!&\!-q_{\phi}\mathcal{A}_{34}^{a}\\ {\mathcal{A}}_{14}\!&\!{\mathcal{A}}_{24}\!&\!\mathcal{A}_{34}\!&\!\mathcal{A}_{44}\end{bmatrix}\!\!,$ (13)
where the new entries $\mathcal{A}^{a}_{ij}$ and $\mathcal{A}^{b}_{ij}$ can be found in (24) in the Appendix. In particular, the new transformation $\phi=q_{\phi}(1-\zeta_{\phi})$ relaxes the condition in (15) by eliminating cross terms arising from the multiplication of $\mathcal{A}_{3}$ with $\mathcal{A}_{2}$ and $\mathcal{A}_{1}$, given that $\mathds{E}[\zeta_{\phi}]=0$ (cf. proof of Theorem 1). It also facilitates the treatment of bilinearity, as will be discussed in Subsection III-B, enabling us to derive the linear conditions presented in (17). With the augmented networked control system formulated, we now introduce the CBC for the dt-ASLS-$\Lambda$ in (11) in the next section.
@@ III Control Barrier Certificate
Consider a dt-SLS-$\Sigma=(X,U,A,B,w,\tau,\theta,\phi)$ in Definition 1, with the safety specification $\Upsilon=(X_{0},X_{1},\mathcal{T})$, and $\hat{x}_{k}^{c}$, $\hat{u}_{k}^{c}$ as the state and input of the controller in (3). A function $\mathbb{B}\!:X\!\times\!X\!\times\!U\!\times\!U\to\mathbb{R}^{+}_{0}$ is said to be a control barrier certificate (CBC) for the augmented dt-ASLS-$\Lambda$ if there exist $\beta,\eta\in\mathbb{R}^{+}_{0}$ with $\beta>\eta$, and $c\in\mathbb{R}^{+}$, such that
@@
Remark 3 (On the multiple unsafe regions)
In the case of multiple unsafe regions, condition (14b) should be enforced separately for each region (cf. simulation results).
@@
The CBC condition (14c) is influenced by the uplink delay $\tau$ and probabilistic data loss in both uplink ($\theta$) and downlink ($\phi$) channels, as these factors explicitly appear in the dt-ASLS-$\Lambda$ described in (10). Consequently, these communication issues directly affect the CBC $\mathbb{B}$ and, in turn, influence the values of the level sets $\eta$ (14a), $\beta$ (14b), and the constant $c$. These parameters collectively quantify the lower bound on the safety probability $1-\xi_{\tau\theta\phi}$, as outlined in Theorem 2.
@@ III-A Computation of CBC and Safety Controller
We consider the structure of our CBC to be quadratic as $\mathbb{B}(\mathcal{Z})=\mathcal{Z}^{\top}\mathcal{P}\mathcal{Z}$, where $\mathcal{P}\in\mathbb{R}^{\kappa\times\kappa}$, with $\kappa=2\varpi+\psi+2n+m$$\mathcal{P}\succ 0$. We then propose a framework for designing a CBC and its corresponding safety controller to enforce safety requirements for the dt-SLS-$\Sigma$. In our approach, we address the first two conditions in Definition 3 by deriving their closed-form expressions, while the third condition is reformulated as a matrix inequality (cf. Assumption 1). To achieve this, we first raise the following assumption.
@@
The maximum delay and data loss that a control system can tolerate depend majorly on the eigenvalues of the state matrix $A$. This is reflected in the matrix inequality (15). Generally, state matrices with larger eigenvalues severely limit the tolerable delay and data loss, whereas state matrices with eigenvalues placed well inside the unit circle allow for larger delays and higher data loss while fulfilling Assumption 1, [29, 30].
@@ III-B Handling Bilinearity
Before presenting the main results of this work under Assumption 1, we note that there is a bilinearity between $\mathcal{P}$ and $\mathcal{F}$ in the matrix inequality (15). To resolve it, we define weight coefficients $\delta_{i}\in[0,1],i\in\{1,2,3,4,5\},$ such that $\sum_{i=1}^{5}\delta_{i}=1$, and
@@
4:     if Matrix inequalities in (17a)–(17d) are solved using $\delta$ and $\mathcal{P}$ in step 3 then
@@
Consider the dt-SLS-$\Sigma=(X,U,A,B,w,\tau,\theta,\phi)$ described in Definition 1. Let Assumption 1 hold, and suppose there exist $\beta,\eta\in\mathbb{R}^{+}_{0}$ with $\beta>\eta$ such that
@@
Consider a dt-SLS-$\Sigma=(X,U,A,B,w,\tau,\theta,\phi)$ as in Definition 1 subject to delay and data loss. Let a CBC $\mathbb{B}$ exist for the augmented system dt-ASLS-$\Lambda$ in (11) as in Definition 3. Then the probability that a trajectory $\mathsf{x}_{k}$ of dt-SLS-$\Sigma$, starting from any initial state $x_{0}\in X_{0}$ under an input signal $u_{k}$ and process noise $w_{k}$ reaches $X_{1}$ within a time horizon $k\in[0,\mathcal{T}]$ is
@@
Then the proposed bound in (21) can be obtained by applying [33, Theorem 3] to (21) and employing conditions (14a) and (14c) in Definition 3. ∎
@@
Algorithm 2 CBC and Safety Controller Computation

### v2

@@ II-C Augmented dt-SLS-$\Sigma$
To formally analyze safety of the networked control system dt-SLS-$\Sigma$, we first introduce an augmented networked system representation that captures both the inputs and states of the dt-SLS-$\Sigma$ and serves as the foundation for defining the CBC, which follows next. The dynamics of $x_{k}$ and $\hat{x}_{k}^{c}$ are given in (3) and (4), while dynamics of $u_{k}$ and $\hat{u}_{k}^{c}$ can be defined as
@@
where $\mathcal{A}\in\mathbb{R}^{\kappa\times\kappa}$ and $\mathcal{D}\in\mathbb{R}^{\kappa\times\psi}$, with its equivalent representation
$\!\!\Lambda\!$ $:\!\!\begin{bmatrix}{\mathbf{x}}_{k+1}\\ \hat{\mathbf{x}}_{k+1}^{c}\\ {\mathbf{u}}_{k+1}\\ \hat{\mathbf{u}}_{k+1}^{c}\end{bmatrix}\!\!=\!\!\underbrace{\begin{bmatrix}\mathcal{A}_{11}&\mathcal{A}_{12}&\mathcal{A}_{13}&\mathcal{A}_{14}\\ \mathcal{A}_{21}&\mathcal{A}_{22}&\mathcal{A}_{23}&\mathcal{A}_{24}\\ \mathcal{A}_{31}&\mathcal{A}_{32}&\mathcal{A}_{33}&\mathcal{A}_{34}\\ \mathcal{A}_{41}&\mathcal{A}_{42}&\mathcal{A}_{43}&\mathcal{A}_{44}\end{bmatrix}}_{\mathcal{A}}\!\begin{bmatrix}{\mathbf{x}}_{k}\\ \hat{\mathbf{x}}_{k}^{c}\\ {\mathbf{u}}_{k}\\ \hat{\mathbf{u}}_{k}^{c}\end{bmatrix}\!\!+\!\mathcal{D}\mathbf{w}_{k},$ (10)
$\text{with}~\mathcal{D}=\!\begin{bmatrix}\mathbf{I}_{n}&{A}&\dots&{A}^{\tau}\\ \boldsymbol{0}_{(\kappa-n)\times n}&\boldsymbol{0}_{(\kappa-n)\times n}&\dots&\boldsymbol{0}_{(\kappa-n)\times n}\end{bmatrix}_{\kappa\times\psi}\!\!\!\!\!\!\!\!\!\!\!\!,$
where $\mathcal{A}_{ij},~i,j\in\{1,\dots,4\}$ are defined as in (23) in the Appendix. Since $\phi_{k-t-1}$ is an i.i.d. Bernoulli random variable for all $t\in\{0,1,\dots,\tau-1\}$, we simplify the notation by denoting $\phi_{k-t-1}$ as $\phi$ and $1-\phi_{k-t-1}$ as $\bar{\phi}$. We used the padding technique [19] in constructing matrix $\mathcal{A}$ to explicitly represent the state variable transitions from one time instant to the next. The padding incorporates trivial equations $x_{i}=x_{i},\,\hat{u}^{c}_{i}=\hat{u}^{c}_{i}$, with $\forall i\in[k-\tau+1,k]$, $\hat{x}_{k}^{c}=\hat{x}_{k}^{c}$, and ${u}_{i}={u}_{i}$, with $\forall i\in[k-\tau,k]$ ensuring continuity in the state representation of the augmented system dt-ASLS-$\Lambda$.
We now decompose matrix $\mathcal{A}$ in (10) into a sum of multiple components. Specifically, $\mathcal{A}$ is expressed as a combination of a baseline matrix $\mathcal{A}_{1}$ and additional terms that depend on the Bernoulli-distributed variable $\phi$. This decomposition allows us to analyze the dynamics of the augmented system dt-SLS-$\Lambda$ by isolating the effect of $\phi$. Furthermore, we use the transformation $\phi=q_{\phi}(1-\zeta_{\phi})$, where the new variable $\zeta_{\phi}$ has a two-point distribution and takes one of two possible values: either $1$ or $1-\frac{1}{q_{\phi}}$, with $\mathds{E}[\zeta_{\phi}]=0$ and $\mathsf{Var}[\zeta_{\phi}]=\mathds{E}[\zeta_{\phi}^{2}]=\frac{1-q_{\phi}}{q_{\phi}}$. This transformation simplifies the computation by allowing the decomposition of $\mathcal{A}$ as
@@
$\mathcal{A}_{1}$ $\!\!=\!\!\begin{bmatrix}\mathcal{A}_{11}\!\!&\!\!\mathcal{A}_{12}\!\!&\!\!\mathcal{A}_{13}^{a}\!\!&\!\!\mathbf{0}_{\psi\times\varpi}\\ {\mathcal{A}}_{21}\!\!&\!\!{\mathcal{A}}_{22}\!\!&\!\!\mathcal{A}_{23}\!\!&\!\!\mathcal{A}_{24}\\ \!\mathbf{0}_{(\varpi+m)\times\psi}\!\!&\!\!\mathbf{0}_{(\varpi+m)\times 2n}\!\!&\!\!\mathcal{A}^{a}_{33}\!\!&\!\!\mathbf{0}_{(\varpi+m)\times\varpi}\!\!\!\\ \mathbf{0}_{\varpi\times\psi}\!\!&\!\!\mathbf{0}_{\varpi\times 2n}\!\!&\!\!\mathbf{0}_{\varpi\times(\varpi+m)}\!\!&\!\!\mathbf{0}_{\varpi\times\varpi}\end{bmatrix}\!\!$,
$\mathcal{A}_{2}$ $\!\!=\!\!\begin{bmatrix}\mathbf{0}_{\psi\times\psi}\!&\!\mathbf{0}_{\psi\times 2n}\!&\!(1-q_{\phi})\mathcal{A}_{13}^{b}\!&\!q_{\phi}\mathcal{A}_{14}^{a}\\ \mathbf{0}_{2n\times\psi}\!&\!\mathbf{0}_{2n\times 2n}\!&\!\mathbf{0}_{2n\times(\varpi+m)}\!&\!\mathbf{0}_{2n\times\varpi}\\ q_{\phi}\mathcal{A}_{31}^{a}\!&\!q_{\phi}\mathcal{A}_{32}^{a}\!&\!(1-q_{\phi})\mathcal{A}_{33}^{b}\!&\!q_{\phi}\mathcal{A}_{34}^{a}\\ {\mathcal{A}}_{41}\!&\!{\mathcal{A}}_{42}\!&\!\mathcal{A}_{43}\!&\!\mathcal{A}_{44}\end{bmatrix}\!\!$,
$\mathcal{A}_{3}$ $\!\!=\!\!\begin{bmatrix}\mathbf{0}_{\psi\times\psi}\!&\!\mathbf{0}_{\psi\times 2n}\!&\!q_{\phi}\mathcal{A}_{13}^{b}\!&\!-q_{\phi}\mathcal{A}_{14}^{a}\\ \mathbf{0}_{2n\times\psi}\!&\!\mathbf{0}_{2n\times 2n}\!&\!\mathbf{0}_{2n\times(\varpi+m)}\!&\!\mathbf{0}_{2n\times\varpi}\\ -q_{\phi}\mathcal{A}_{31}^{a}\!&\!-q_{\phi}\mathcal{A}_{32}^{a}\!&\!q_{\phi}\mathcal{A}_{33}^{b}\!&\!-q_{\phi}\mathcal{A}_{34}^{a}\\ \mathbf{0}_{\varpi\times\psi}\!&\!\mathbf{0}_{\varpi\times 2n}\!&\!\mathbf{0}_{\varpi\times(\varpi+m)}\!&\!\mathbf{0}_{\varpi\times\varpi}\end{bmatrix}\!\!$, (12)
and the new entries $\mathcal{A}^{a}_{ij}$ and $\mathcal{A}^{b}_{ij}$ can be found in (24) in the Appendix. In particular, the new transformation $\phi=q_{\phi}(1-\zeta_{\phi})$ relaxes the condition in (14) by eliminating cross terms arising from the multiplication of $\mathcal{A}_{3}$ with $\mathcal{A}_{2}$ and $\mathcal{A}_{1}$, given that $\mathds{E}[\zeta_{\phi}]=0$ (cf. proof of Theorem 1). It also facilitates the treatment of bilinearity, as will be discussed in Subsection III-B, enabling us to derive the linear conditions presented in (16). With the augmented networked control system formulated, we now introduce the CBC for the dt-ASLS-$\Lambda$ in (10) in the next section.
@@ III Control Barrier Certificate
Consider a dt-SLS-$\Sigma=(X,U,A,B,w,\tau,\theta,\phi)$ in Definition 1, with the safety specification $\Upsilon=(X_{0},X_{1},\mathcal{T})$, and $\hat{x}_{k}^{c}$, $\hat{u}_{k}^{c}$ as the state and input of the controller in (3). A function $\mathbb{B}\!:X\!\times\!X\!\times\!U\!\times\!U\to\mathbb{R}^{+}_{0}$ is said to be a control barrier certificate (CBC) for the augmented dt-ASLS-$\Lambda$ if there exist $c,\beta,\eta\in\mathbb{R}^{+}$ with $\beta>\eta$, such that
@@
The CBC condition (13c) is influenced by the uplink delay $\tau$ and probabilistic data loss in both uplink ($\theta$) and downlink ($\phi$) channels, as these factors explicitly appear in the dt-ASLS-$\Lambda$ described in (9). Consequently, these communication issues directly affect the CBC $\mathbb{B}$ and inevitably influence the values of the level sets $\eta$ (13a), $\beta$ (13b), and the constant $c$. These parameters collectively quantify the lower bound on the safety probability $1-\xi_{\tau\theta\phi}$, as outlined in Theorem 2.
Remark 4 (On downlink delays)
Our framework can be readily extended to handle downlink delays using a similar approach as for the uplink. However, for clarity of presentations, we leave this extension to future works.
@@ III-A Computation of CBC and Safety Controller
We consider the structure of our CBC to be quadratic as $\mathbb{B}(\mathcal{Z})=\mathcal{Z}^{\top}\mathcal{P}\mathcal{Z}$, where $\mathcal{P}\in\mathbb{R}^{\kappa\times\kappa}$, with $\kappa=2\varpi+\psi+2n+m$, and $\mathcal{P}\succ 0$. We then propose a framework for designing a CBC and its corresponding safety controller to enforce safety requirements for the dt-SLS-$\Sigma$. To achieve this, we first raise the following assumption.
@@
The maximum delay and data loss that a control system can tolerate depend majorly on the eigenvalues of the state matrix $A$, which is reflected in the matrix inequality (14). Generally, state matrices with larger eigenvalues severely limit the tolerable delay and data loss, whereas state matrices with eigenvalues placed well inside the unit circle allow for larger delays and higher data loss while fulfilling Assumption 1, [23, 24].
@@ III-B Handling Bilinearity
Before presenting the main results of this work under Assumption 1, we note that there is a bilinearity between $\mathcal{P}$ and $\mathcal{F}$ in the matrix inequality (14). To resolve it, we define weight coefficients $\delta_{i}\in(0,1),i\in\{1,2,3,4,5\},$ such that $\sum_{i=1}^{5}\delta_{i}=1$, and
@@
4:  if Matrix inequalities in (16a)–(16d) are solved using $\delta$ and $\mathcal{P}$ in step 2 then
@@
Consider the dt-SLS-$\Sigma=(X,U,A,B,w,\tau,\theta,\phi)$ described in Definition 1. Let Assumption 1 hold, and suppose there exist $\beta,\eta\in\mathbb{R}^{+}$ with $\beta>\eta$ such that
@@
Consider a dt-SLS-$\Sigma=(X,U,A,B,w,\tau,\theta,\phi)$ as in Definition 1 subject to delay and data loss. Let there exist a CBC $\mathbb{B}$ for the augmented system dt-ASLS-$\Lambda$ in (10) as in Definition 3. Then the probability that a trajectory $\mathsf{x}_{k}$ of dt-SLS-$\Sigma$, starting from any initial state $x_{0}\in X_{0}$ under an input signal $u_{k}$ and process noise $w_{k}$ reaches $X_{1}$ within a time horizon $k\in[0,\mathcal{T}]$ is
@@
Then the proposed bound in (20) can be obtained by applying [27, Theorem 3] to (20) and employing conditions (13a) and (13c) in Definition 3. ∎
@@
The procedure for constructing the CBC and its controller under communication constraints is outlined in Algorithm 2.
Algorithm 2 CBC and safety controller construction